Jim, Carla and Tommy are all from one family. Carla is 5 years older than Jim. Tommy is 6 years older than Carla. How old is Tommy?
Answer by mananth(16946)

Jim, Carla and Tommy are all from one family. Carla is 5 years older than Jim. Tommy is 6 years older than Carla. How old is Tommy?
...
let Jim be x years
carla = x+5
Tommy = x+5+6=x+11 years
...